A fountain shoots water from a nozzle at its base. The height h, in feet, of the water above the ground t seconds after it leaves the nozzle is given by the function h(t) = –16t2 + 17t + 8. What is the maximum height of the water? Round to the nearest tenth
Answer by ewatrrr(24785)

h(t) = –16t2 + 17t + 8 = -16(t-17/32)^2 + 17^2/64 + 512/64
max = 801/64 = 12.5ft (nearest tenth)